A firm depreciated its assets with regard to an investment to $0.00 book value by the end of the project's 15 year life.   However, the firm expects to sale the assets to have a scrap market value of $1,200,000 in 15 years. What is the PV of the after-tax salvage value if r = 14% per year, and the firm's tax rate is 28%?
Answer #1

Step 1: Find after-tax cash flow:

\textup{After-tax cash flow} = \textup{Cash flow }\times (1-\textup{tax rate})

= \textup{1,200,000 }\times (1-\textup{28\%})

= \textup{1,200,000}\times 72\%

= \textup{\$\,864,000 }

Step 2: Find the present value of after tax cash flow:

\textup{PV} = \frac{\textup{CF}}{(1+r)^{n}}

Where,
PV = Present value
CF = Future cash flow
r = discount rate
n = number of years

\therefore \textup{PV} = \frac{\textup{ 864,000 }}{(1+0.14)^{15}}

= \frac{\textup{ 864,000 }}{7.137938}

= \textup{\$\,121,043.36 }

Therefore, PV of the after-tax salvage value is $121,043.36